Accelerator Update for July 18, 2008 - July 21, 2008
Friday July 18
The day shift began with Operations monitoring the Tevatron (TeV) store 6302 (41.56E30) and the Recycler (RR) stash (96.4E10), with the Antiproton Source stacking (12.3E10), with the Main Injector, Booster, and Linac operating normally, with the Collider Detector at Fermilab (CDF) and DZero (D0) using the colliding beams, with the SciBooNE / MiniBooNE (BNB) experiments off, and with the NuMI experiment taking beam.
Operations began trying to send beam to BNB at 8:25 AM.
Operations terminated store 6302 at 8:36 AM.
Operations reported at 9:02 AM that they were having problems getting some of the quadrupole power supplies on for BNB.
Cryo system experts reported at 9:07 AM that they were going to work on the MYCOM compressors at TeV sectors B0 and BA.
At 9:26 AM, Operators responded to many reports of a bad smell in the Booster West parking lot. They found that the smell was coming for the West Anode Power Supply (WAPS) and the Booster West RF stations. Operations contacted an RF expert.
Operations put Pbar in standby at 9:42 AM due to the WAPS problem.
At 9:53 AM, Operations reported that the BNB turn on problems were due to the quadrupoles and MI-10. An expert turned them on.
Cryo system experts reported at 9:54 AM that the sector E4 wet engine bearing failure would take about eight-hours to repair. This work won’t begin until after the TeV study ends.
Operations reported at 10:05 AM that a WAPS resistor had failed and that repairs were underway.
The TeV studies ended at 10:33 AM.
Operations prepared D0 for an access at 10:40 AM.
The Booster WAPS repair was completed by 10:55 AM.
At 11:01 AM, Operations reported that BRF15 had tripped off twice. An RF expert investigated.
Operations prepared CDF for an access at 11:12 AM.
At 12:01 PM, a Linac Klystron RF station (KRF1) tripped off and would not reset. A Linac expert reported that its PFN safety relay was not working and he was beginning repairs.
Operations put Pbar into standby at 12:20 PM.
A LCW expert reported at 1:09 PM that he brought column #2 online at CUB.
Experts reported around 1:45 PM that the TeV sector C2 cryo system that they had a problem with a valve (EVBY) and that the magnets had warmed up to ~19 Kelvin. Operations contacted vacuum experts to check the C2 vacuum.
CDF and D0 came out of access at 2:13 PM.
At 2:23 PM, a Linac expert reported that the he had replaced the PFN safety relay for KRF1, but the replacement relay was not working. The investigation continued.
At 2:51 PM, KRF5 tripped off. It reset.
Operations reported at 2:53 PM that KRF1 was back online. Linac resumed sending beam.
At 2:59 PM, BRF15 tripped off. It reset.
Operations reported at 4:05 PM that the Recycler had stashed antiprotons three times during the day shift.
At 4:25 PM, BRF3 tripped off for the first of two times in 32 minutes. It reset.
At 7:02 PM, BRF15 tripped off. It reset.
Operations established beam to MTest at 8:06 PM.
Operations put the TeV through a dry squeeze at 9:20 PM.
At 10:10 PM, Operations reported that the Linac Klystron Vernier had tripped off three times in 10 minutes on a LCW flow fault. It reset.
Operations put the TeV through a wet squeeze at 10:13 PM.
Operations put the TeV into shot setup at 10:54 PM.
At 11:04 PM, BRF3 tripped off. It reset.
Operations reported at midnight that the Recycler had stashed four times during the evening shift.
Saturday July 19
The midnight shift began with the TeV in shot setup, with Operations monitoring the stash (9.9E10), with Pbar stacking (40.6E10), and with BNB and MTest taking beam.
Operations established store 6305 at 12:29 AM with an initial luminosity of 238.29E30.
At 1:22 AM, the MI Power Supply interlock Loop tripped off on a ground fault indication. It reset, but then tripped off again. Operations conducted a hipot, found a fault and then contacted an expert.
Operations reported at 2:07 AM that while talking with a MI expert, the MI fault cleared. Pbar resumed stacking.
At 8:03 AM BRF14 tripped off. It would not reset. Operations contacted an RF expert.
Operations reported at 8:09 AM that the Recycler had stashed six times during the midnight shift.
At 8:13 AM, BRF14 was bypassed. Operations lowered the Booster beam intensity due to the number of RF stations that were offline.
At 9:06 AM, Operations reported that the Klystron Debuncher gradient was falling out and causing Booster trips. Linac experts were called in.
At 9:31 AM, a Pbar Accumulator quadrupole (A:LQ) power supply tripped off. The stack was lost, which was 38.0E10. Operations called in an expert.
At 9:56 AM, BRF3 tripped off. It reset.
Operations disabled beam to MTest at 10:10 AM.
A power supply expert reported that he checked out the Accumulator A:LQ power supply and found nothing wrong. Pbar resumed stacking at 10:35 AM.
At 1:03 PM, BRF15 tripped off for the first of two times in one hour. It reset.
At 6:43 PM, LRF4 tripped off. It reset.
At 7:14 PM, four Pbar Debuncher beam valves closed. There was no vacuum problem and the beam valves opened.
Operations reported at 7:58 PM that the Recycler had stashed six times during the last 12 hours.
Operations terminated store 6305 at 8:04 PM.
Operations put the TeV into shot setup at 8:32 PM.
At 9:10 PM, MI suffered a ramp dump on a ground fault indication. Everything reset.
Operations established store 6306 at 9:55 PM with an initial luminosity of 290.09E30.
At 10:25 PM, KRF2 tripped off. It reset.
At 11:03 PM, the Linac Klystron Vernier tripped off on a LCW flow fault. It reset.
At 11:21 PM, the NuMI beam permit tripped off. It reset.
Sunday July 20
The midnight shift began with Operations monitoring store 6306 and the stash, with Pbar stacking, and with BNB, NuMI, and the MTest experiments taking beam.
At 1:31 AM, LRF4 tripped off. It reset.
At 2:06 AM, MIRF18 tripped off and took down the NuMI beam permit. It reset, but continued to trip.
At 3:33 AM, Operations bypassed MIRF18.
At 6:41 AM, MIRF5 tripped off and would not reset. Operations called in an expert.
Operations reported at 7:54 AM that the Recycler had stashed seven times during the last 12 hours.
Operations reported at 8:48 AM that a RF expert had brought MIRF5 back online. He changed out a power supply.
Operations reported at 9:22 AM that MIRF18 was back online. The expert reported finding a mod block problem.
At 11:05 AM, KRF3 tripped off. It reset.
At 11:35 AM, KRF5 tripped off for the first of five times over the next five hours. It reset.
Operations disabled the MTest beam at 1:01 PM.
At 1:33 PM, Operations reported that MIRF18 began tripping off again. Operations contacted experts.
Operations terminated store 6306 at 3:06 PM.
Operations put the TeV into shot setup at 3:17 PM.
At 4:22 PM, the MI PSLOOP tripped off and would not reset.
At 4:49 PM, Operations reported finding a MI ground fault with a hipot. Operations contacted an expert. The fault cleared, but the expert thought the problem was with an MI-10 power supply.
Operations put the TeV into shot setup at 5:28 PM.
Operations established store 6308 at 7:09 PM with an initial luminosity of 291.15E30.
Operations established beam to MTest at 7:37 PM.
Operations reported at 8 PM that the Recycler had stashed six times during the last 12 hours.
At 10:04 PM, the MI PSLOOP tripped off. It reset.
Monday July 21
The midnight shift began with Operations monitoring store 6308 and the stash, with Pbar stacking, and with BNB, NuMI, and the MTest experiments taking beam.
At 4:13 AM, the Klystron Vernier tripped off. It reset.
Operations terminated store 6308 around 5:44 AM.
Operations prepared CDF and D0 for access at 6 AM.
Operations prepared the TeV transfer hall for access at 7:28 AM.
Operations prepared Pbar for access at 7:40 AM.
Operations reported at 7:58 AM that the Recycler had stashed eight times during the last 12 hours.
The Plans for Monday
The plans for today are to locate the TeV ground fault and make necessary repairs and upgrades in Pbar. The other machines will make use of this time and make necessary repairs as needed. D0 and CDF accessed their collision halls.
Comments and Suggestions
|last modified 07/21/2008 email Fermilab|
|Security, Privacy, Legal|